# Behind the Magic, MERLIM: Multi-modal Evaluation Benchmark for Large Image-Language Models

### v2

@@ A.3 LLMs vs IT-LVLMs
A.4 Gradient Analysis
To further investigate the visual grounding of IT-LVLMs, we take follow(Selvaraju et al., 2017) and compute the proportion of the total gradient attributable to each specific type of input (Image, Question, Answer) for each predicted token. We analyze the model’s output logits before token selection and propagate the gradient to the input by identifying the maximum activation value per output token. Specifically, for each output token, we determine the gradient contribution from each input type and then average these contributions across all answers. To simplify our analysis, we focus on architectures with autoregressive LLMs, such as Vicuna, which explicitly use previous output tokens as inputs for predicting subsequent tokens.
@@ A.4 Gradient Analysis
As shown in Table 5, the gradient of the visual input is significantly lower than that of the language inputs (Question and Answer). This indicates that LLMs prioritize language tokens over visual ones when predicting answers, leading to hallucinations based on language biases. Additionally, we observe that only a few tokens account for most of the visual gradients. For instance, in InstructBLIP with Vicuna-7B, just 10 out of 30 tokens represent nearly 73% of the visual gradients, making it challenging for specific and strong visual information to be adequately considered. These findings support the results of our previous experiments.
@@
A.5 Number of unsuitable outputs

### v3

@@ 8 LLMs vs IT-LVLMs
9 Gradient Analysis
@@
To further investigate the visual grounding of IT-LVLMs, we take inspiration from [47] and compute the proportion of the total gradient attributable to each specific type of input (Image, Question, Answer) for each predicted token. We analyze the model’s output logits before token selection and propagate the gradient to the input by identifying the maximum activation value per output token. Specifically, for each output token, we determine the gradient contribution from each input type and then average these contributions across all answers. To simplify our analysis, we focus on architectures with autoregressive LLMs, such as Vicuna, which explicitly use previous output tokens as inputs for predicting subsequent tokens.
As shown in Figure 5, the gradient of the visual input is significantly lower than that of the language inputs (Question and Answer). This indicates that LLMs prioritize language tokens over visual ones when predicting answers, leading to hallucinations based on language biases. Additionally, we observe that only a few tokens account for most of the visual gradients. For instance, in InstructBLIP with Vicuna-7B, just 10 out of 30 tokens represent nearly 73% of the visual gradients, making it challenging for specific and strong visual information to be adequately considered. These findings support the results of our previous experiments.
@@
10 Number of unsuitable outputs
@@
11 Visual Examples of MERLIM’s tasks.
Figure 10 presents additional visualizations of the tasks evaluated by MERLIM. It shows the InstructBLIP outputs for both the original and edited images across the three tasks. Notably, the model consistently produces visually ungrounded predictions (hallucinations) in all scenarios, highlighting the complexity of the hallucination problem in these instructional models.
Figure 10: Visual Examples. We present additional visual examples illustrating the tasks evaluated by MERLIM. Sub-figures (a), (b), and (c) showcase comparisons of InstructBLIP outputs for original and edited images, focusing on the tasks of Object Recognition, Inter-Object Relationship Understanding, and Object Counting, respectively.
12 Impact Statements
MERLIM is a benchmark designed to assess the performance of Instruction Tuning Large Vision and Language Models (IT-LVLMs). Besides the empirical evaluation, MERLIM’s primary aim is to identify and quantify instances of "Hallucination" events in the textual responses generated by IT-LVLMS. Consequently, MERLIM represents a tool with a potentially positive societal impact by encouraging advanced IT-LVLM models to be more robust to hallucination events and, therefore, bring more factual and informative responses.